# Constraints on the Hubble and matter density parameters with and without modelling the CMB anisotropies

### v1

@@ Abstract
journal: The Open Journal of Astrophysics
@@ 2.2 The horizon scale at matter-radiation equality
Surveys of large-scale structure in redshift space provide the tight constraint $k_{\mathrm{eq}}/h=\left(1.64\pm 0.05\right)\times 10^{-2}$/Mpc (Philcox et al., 2022). This is important to our discussion because $k_{\mathrm{eq}}/h\propto\Omega_{\mathrm{M}}h$ (equation 3 of Eisenstein & Hu, 1998, note they use $\Omega_{0}$ to denote $\Omega_{\mathrm{M}}$). We obtain the proportionality constant by noting that Philcox et al. (2022) report most likely values of $h=0.648$ and $\Omega_{\mathrm{M}}=0.338$ (see their figure 2). This combination presumably corresponds to their most likely $k_{\mathrm{eq}}/h$. We assume its fractional uncertainty is the same as that in $\Omega_{\mathrm{M}}h$.
@@ 3 Results and Discussion
Alternatively, a modification to the expansion history at late times may have little impact on $t_{\mathrm{U}}$ and the angular diameter distance to the CMB. If the modification is carefully tuned, it could enhance the present expansion rate enough to solve the Hubble tension while not much affecting the other observables. Such a modification may arise from dark energy not being a pure cosmological constant (e.g., Harko, 2023; Yao & Meng, 2023). The main problem with this is that a larger $H_{0}$ requires a higher energy density according to the Friedmann equations, which would mean that the dark energy density rises with time (Dahmani et al., 2023). This leads to a phantom equation of state, which suffers from theoretical issues with vacuum stability (Ludwick, 2017). It would also be unusual if we happened to be living at just the time when dark energy starts deviating from a cosmological constant. Moreover, we would have to neglect the evidence that structure formation is more efficient than expected in $\Lambda$CDM, e.g. from the high mass, redshift, and collision velocity of the El Gordo interacting galaxy clusters (Asencio et al., 2021, 2023). A faster background expansion rate would create more Hubble drag on growing structures, making it more difficult to explain the fast observed bulk flows and form large and deep voids. Even so, a modification to the expansion history at late times would avoid the serious difficulties pointed out in this contribution with modifications at early times, even if such modifications can achieve a good fit to the CMB anisotropies (see also Vagnozzi, 2023, and references therein).
@@ Appendix A The expansion history around matter-radiation equality
where $a$ is the cosmic scale factor with value 1 today, an overdot denotes a time derivative, and $\Omega_{\mathrm{M}}$ ($\Omega_{\mathrm{r}}$) is the present fraction of the cosmic critical density in matter (radiation). We can solve this by separating the variables and using the substitution $a=a_{\mathrm{eq}}/\tan^{2}\theta$. Applying the usual boundary condition that $a=0$ when $t=0$ gives the following result:
@@
Our derivation remains valid if the argument of the central $\operatorname{acos}$ function becomes $>1$, but then we need to use the generalised definition of $\cos$ and $\operatorname{acos}$ for complex numbers. This is equivalent to setting $\cos\to\cosh$ and $\operatorname{acos}\to\operatorname{acosh}$ in Equation A6 and using only real quantities. We can alternatively use this equation as shown, in which case the result of the $\operatorname{acos}$ operation smoothly transitions from the real axis to the imaginary axis of the complex plane, but the $\cos$ operation then returns us back to the real axis.

### v2

@@ Abstract
journal: Astronomy
@@ 2.2 The horizon scale at matter-radiation equality
Surveys of large-scale structure in redshift space provide the tight constraint $k_{\mathrm{eq}}/h=\left(1.64\pm 0.05\right)\times 10^{-2}$/Mpc (Philcox et al., 2022). This is important to our discussion because $k_{\mathrm{eq}}/h\propto\Omega_{\mathrm{M}}h$ (equation 3 of Eisenstein & Hu, 1998, note they use $\Omega_{0}$ to denote $\Omega_{\mathrm{M}}$). We obtain the proportionality constant by noting that Philcox et al. (2022) report most likely values of $h=0.648$ and $\Omega_{\mathrm{M}}=0.338$ (see their figure 2). This combination presumably corresponds to their most likely $k_{\mathrm{eq}}/h$. We assume its fractional uncertainty is the same as that in $\Omega_{\mathrm{M}}h$. We use their study only to constrain this product rather than follow their approach of breaking the degeneracy using a prior on $\Omega_{\mathrm{M}}$, since this often comes from the same information already considered in Section 2.1, which we need to avoid double-counting.
@@ 3 Results and Discussion
An interesting aspect of Figure 1 is that without the age constraint shown in blue, there is excellent agreement between the local $cz^{\prime}$ and the constraints from UCS and $k_{\mathrm{eq}}/h$. This suggests an early time resolution to the Hubble tension that allows the CMB anisotropies to be fit at higher $h$. The problem is the absolute ages of old stars (Cimatti & Moresco, 2023) and the Galactic disc as a whole (Xiang et al., 2024). It is also possible to constrain cosmology using relative stellar ages by treating stellar populations as cosmic chronometers (CCs; Moresco et al., 2018, 2020). The idea of the CC technique is that a ‘red and dead’ galaxy formed its stars in a rapid burst at very early times, with the stellar population passively evolving ever since. Since stars of different masses have different lifetimes, we may observe stellar populations in such galaxies at different redshifts to determine the relative age between those redshifts. This is similar to observing the main sequence turnoff in a globular cluster, but using relative rather than absolute ages. This can make the results less sensitive to evolutionary phases that might be hard to model, provided the model at least captures how the duration of such phases changes with stellar mass. The use of a much larger number of stars also reduces sensitivity to individual stars. The results obtained with CCs are almost entirely from $z>0.3$, making them largely insensitive to the very low redshifts at which a local void might play a role. This allows the CC method to constrain the background cosmology in a somewhat different way to nearby stars. The slope of the age-redshift relation can only constrain some combination of $H_{0}$ and $\Omega_{\mathrm{M}}$, so Cogato et al. (2024) combined CCs with uncalibrated SNe, BAOs, and $\gamma$-ray bursts (GRBs) treated as UCS, thereby obtaining $H_{0}=67.2^{+3.4}_{-3.2}$ km/s/Mpc. This is much more consistent with the Planck value than the local $cz^{\prime}$, which is $\approx 2\sigma$ higher. The good agreement between the constraints from CCs and from the absolute ages of ancient Galactic stars provides additional confidence that $H_{0}$ is below the local $cz^{\prime}$, at least if we assume an FLRW cosmology which is almost homogeneous today on the scales involved in the local $cz^{\prime}$ measurement.
It is therefore difficult to solve the Hubble tension through modifications to the physics at very early times, even if a good fit to the CMB can be retained (a non-trivial task; see Vagnozzi, 2021). Thus, the most promising way to retain the usual assumption that $cz^{\prime}=H_{0}$ is to modify the expansion history at late times. This may have little impact on $t_{\mathrm{U}}$ and the angular diameter distance to the CMB. If the modification is carefully tuned, it could enhance the present expansion rate enough to solve the Hubble tension while not much affecting the other observables (Tiwari et al., 2024). Such a modification may arise from dark energy not being a pure cosmological constant (e.g., Harko, 2023; Yao & Meng, 2023; Rezazadeh et al., 2024). The main problem with this is that a larger $H_{0}$ requires a higher energy density according to the Friedmann equations, which would mean that the dark energy density rises with time (Dahmani et al., 2023). This leads to a phantom equation of state, which suffers from theoretical issues with vacuum stability (Ludwick, 2017). It would also be unusual if we happened to be living at just the time when dark energy starts deviating from a cosmological constant. Moreover, we would have to neglect the evidence that structure formation is more efficient than expected in $\Lambda$CDM, e.g. from the high mass, redshift, and collision velocity of the El Gordo interacting galaxy clusters (Asencio et al., 2021, 2023). A faster background expansion rate would create more Hubble drag on growing structures, making it more difficult to explain the fast observed bulk flows and form large and deep voids. Even so, a modification to the expansion history at late times would avoid the serious difficulties pointed out in this contribution with modifications at early times, even if such modifications can achieve a good fit to the CMB anisotropies (see also Vagnozzi, 2023, and references therein).
@@ Appendix A The expansion history around matter-radiation equality
where $a$ is the cosmic scale factor with value 1 today, an overdot denotes a time derivative, and $\Omega_{\mathrm{M}}$ ($\Omega_{\mathrm{r}}$) is the present fraction of the cosmic critical density in matter (radiation). We can solve this by separating the variables and using the substitution $a=a_{\mathrm{eq}}\sinh^{2}\theta$. Applying the usual boundary condition that $a=0$ when $t=0$ gives the following result:
@@
Our derivation remains valid if the argument of the central $\operatorname{acos}$ function becomes $>1$, but then we need to use the generalised definition of $\cos$ and $\operatorname{acos}$ for complex numbers. This is equivalent to setting $\cos\to\cosh$ and $\operatorname{acos}\to\operatorname{acosh}$ in Equation A6 and using only real quantities. If we instead use this equation as shown, the result of the $\operatorname{acos}$ operation smoothly transitions from the real axis to the imaginary axis of the complex plane, but the $\cos$ operation then returns us back to the real axis.